OpenStudy (anonymous):

a city population is growning at a rate of 3.5% per year. What is the doubling time? Bu what factor will the population increase in 50 years

OpenStudy (anonymous):

if 3.5% is the annual rate, your job it to solve \[(1.035)^t=2\] for \(t\)

OpenStudy (anonymous):

takes exactly one step (and a calculator) by the change of base formula \[\large (1.035)^t=2\iff t=\frac{\ln(2)}{\ln(1.035)}\]

OpenStudy (mathmate):

For rough checking purposes, the rule of 72 may be used, which says that the doubling time is roughly 72/increase rate in %. So the doubling time is roughly 72/3.5=20.5 years. In 50 years, it would have doubled 2.5 times, roughly factor=4*sqrt(2)=5.6

OpenStudy (anonymous):

good rule, that rule of 72!

OpenStudy (mathmate):

yes, it's very handy, especially for mental calculations.